OpenStudy (anonymous):

Assume that the playbook contains 13 passing plays and 9 running plays. The coach randomly selects 7 plays from the playbook. What is the probability that the coach selects at least 2 passing plays and at least 2 running plays?

Directrix (directrix):

P(at least 2 PP and at least 2 RP) = [C(13,2)C(9,5) + C(13,3)C(9,4) + C(13,4)C(9,3) + C(13,5)C(9,2)] / C(22,7) = 3172 / 3553 = .893 approx

OpenStudy (anonymous):

Correct